# Mon CHÈRI $\heartsuit$ Adapting Capability Hardware Enhanced RISC with Conditional Capabilities

### v1

@@ Abstract
Up to 10% of memory-safety vulnerabilities in languages like C and C++ stem from uninitialized variables. This work addresses the prevalence and lack of adequate software mitigations for uninitialized memory issues, proposing architectural protections in hardware. Capability-based addressing, such as the University of Cambridge’s CHERI, mitigates many memory defects, including spatial and temporal safety violations at an architectural level. However, current CHERI designs do not handle undefined behavior from uninitialized variables. We extend the CHERI capability model to include “conditional capabilities”, enabling memory-access policies based on prior operations. This allows enforcement of policies that satisfy memory-safety objectives such as “no reads to memory without at least one prior write” (Write-before-Read). We present our architecture extension, compiler support, and a detailed evaluation of our approach using the QEMU full-system simulator and our modified FPGA-based CHERI-RISCV softcore. Our evaluation shows Write-before-Read conditional capabilities are practical, with high detection accuracy while adding a small ($\approx$ 3.5%) overhead to the existing CHERI architecture.
I Introduction
@@
Uninitialized variables, variables that are declared but not assigned a value, are a well-known source of software defects in the C-family of programming languages. In general, all run-time allocated memory in C, C++, and even in modern languages like Rust start out as uninitialized. In this state, the value of the memory is indeterminate and may not even reflect a valid state for the variable type. Attempting to interpret uninitialized memory will cause undefined behavior, which can cause serious security vulnerabilities. In particular, uninitialized memory may expose residual data from previously deallocated data structures. This may inadvertently result in information disclosure, ranging from leaked pointer values [12] to the exposure of sensitive cryptographic keys [73]. Leaked pointer values can be exploited by attackers to bypass address space layout randomization (ASLR) [12], while the use of uninitialized variables can allow arbitrary code execution attacks [47]. Following recent surveys, use-before-initialized conditions account for a sizable 10% of memory-safety vulnerabilities in the wild [35, 7, 75].
@@
Research receiving particular attention are hardware-assisted defenses against memory-safety issues [89]. This is motivated by the need to reduce the overhead of run-time defenses through hardware/software co-designs and by a willingness by processor manufacturers to incorporate hardware mechanisms for software security into their designs [61, 5, 32]. A prominent research example is Capability Hardware Enhanced RISC Instructions (CHERI) [85], a joint research project of SRI International and the University of Cambridge. CHERI extends several existing hardware instruction-set architectures (ISAs) with novel features to enable fine-grained memory protection and software compartmentalization.
@@
In its default configuration, CHERI provides spatial safety through capability-based addressing. While these safety features allow software to preclude many classes of memory-safety defects CHERI’s capabilities do not address uninitialized variables. Microsoft Security Response Center (MSRC) conducted a comprehensive analysis of vulnerabilities reported in 2019 to assess the CHERI ISAv7 [35]. The findings indicate that only 31% of the reported vulnerabilities could have been mitigated through the default configuration of CHERI. An additional 24% could be mitigated by CHERI when configured to provide partial temporal safety under the Cornucopia mechanism [82]. The analysis further highlights that at least 12% of the assessed vulnerabilities could have been mitigated if CHERI would protect against uninitialized access. This work explores and evaluates extensions for CHERI to mitigate those 12%.
@@
In this paper, we extend the CHERI capability model to express memory-access policies that eliminate undefined behavior associated with uninitialized memory. We introduce conditional permissions (CPs) to capability-based addressing. CPs can express memory-access policies that take previous operations on a memory area into account. This enables conditional capabilities with fine-grained policies that satisfy memory-safety objectives such as “no reads of memory which has not been the subject of at least one write” (Write-before-Read) or “this memory can be written to only once” (Write-Once). We describe these CPs in § III. At a high level, CPs are enforced by introducing the notion of operation-specific bounds to capability-based address schemes:
  • We introduce conditional permissions and conditional capabilities for capability-based addressing (§ III).
@@
  • We integrate conditional capabilities to CHERI-RISC-V in prototypes based on the QEMU full-system emulator and an FPGA softcore based on Flute64Cheri IP (§ IV).
@@
  • We add support for Write-before-Read conditional permissions to the CHERI-enabled Clang/LLVM compiler (§ IV-C) and two embedded memory allocators (§ IV-D).
@@
  • We evaluate CP security using $>1000$ test cases for uninitialized variables in the NIST Juliet test suite and performance using EEMBC CoreMark benchmarks (§ V).
@@
Our results show that we achieve 100% detection with only six false positives ($\approx$ 1%) for the test suite, where the false positives do exhibit uninitialized (but non-vulnerable) memory access behavior. We further show that we impose performance overheads of only $\approx$ 3.5% over CHERI (7% combined). In summary, our work effectively and efficiently addresses uninitialized memory vulnerabilities, combining very high detection accuracy with performance penalties that are substantially lower than other detection mechanisms in hardware and software.
II Background
In the following, we report on the prevalence of uninitialized memory vulnerabilities based on statistics from the Common Vulnerability Enumeration (CVE) program (§ II-A) and provide necessary background on hardware-based mitigation against memory-safety vulnerabilities, principally capability-based addressing (§ II-B) and CHERI (§ II-C).
II-A How prevalent are uninitialized memory vulnerabilities?
@@
MSRC reports that, between 2017 and 2019, uninitialized memory vulnerabilities accounted for $\approx$ 5-10% of the CVEs issued by Microsoft[35, 7]. In 2019, uninitialized memory accesses were the fourth largest class of memory defects (10%), after spatial (44%), temporal (29%), and type confusion (14%). Similarly, a more recent survey of memory-safety CVEs between 2015 – 2022 by Herb Sutter [75] indicates that use-before-initialized conditions account for $\approx$ 9% of all reported memory-safety vulnerabilities, after lifetime safety (49%), bounds safety (18%), and type confusion (11%), indicating that uninitialized memory accesses remain an issue in industrial code bases that impacts the operational security of systems.
II-B Capability-Based Addressing
@@
Interest in capability-based addressing diminished with the introduction of memory management units (MMUs) that, in addition to performing address translation between virtual and physical memory addresses, also manage access control to virtual memory. However, capabilities differ fundamentally from the access control provided by MMUs in an important way: whereas MMUs associate permissions to individual memory pages, capabilities associate permissions to the references used to address memory.
II-C Capability Hardware Enhanced RISC Instructions
@@
CHERI, which stands for Capability Hardware Enhanced RISC Instructions, is an ISA extension for a capability-based architectural protection model developed over ten years of DARPA-supported hardware-software co-design. The CHERI architecture extends an underlying conventional ISA with hardware-supported capabilities that are used to protect virtual addresses used as code or data pointers. The CHERI ISA specification [79] defines the representation of capabilities held in registers and memory, as well as capability-aware instructions to manipulate them. Currently, implementations of the CHERI capability architecture exist for MIPS, RISC-V, and ARMv8-A instruction sets. CHERI-enabled processors have been developed by Arm [29], Microsoft [2], as well as companies and universities part of the RISC-V ecosystem.
@@
Figure 1 shows the in-memory representation of a capability. Each CHERI capability is double the width of the native integer pointer type of the baseline architecture: 128 bits on 64-bit platforms and 64 bits on 32-bit platforms. One additional bit, the validity tag ➀, is stored separate from the capability and is used to protect the integrity of the capability: any manipulation of the capability in-memory by non-capability-aware instructions invalidates the tag and its corresponding capability. Capability-aware instructions maintain the tag as long as certain architectural invariants are met. This integrity-protection mechanism prevents direct in-memory manipulations and injection of arbitrary data as capabilities. The permissions ➁ control how the capability can be used and consists of the permissions described in § II-B. The object type ➂ allows capabilities to be temporarily "sealed", which renders the capability unusable until it is "unsealed" by a special instruction. Sealing is used in the CHERI architecture to implement opaque pointer types and fine-grained in-process isolation. The bounds ➃ describe a lower and upper bound relative to the baseline architecture address ➄, which limits the portion of address space the capability is able to access. To reduce the in-memory footprint of capabilities, the bounds are stored in a compressed format [84] with both bounds in 28 bits (for a 64-bit address), loosing precision as the object size increases.
@@
New capabilities in the CHERI architecture are always derived from a pre-existing capability that traces their heritage back to initial capabilities made available to firmware at boot time. CHERI enforces monotonicity on newly created capabilities, ensuring that the new capability constructed by a capability-aware instruction cannot possess permissions or bounds that exceed those of the original capability. The only exceptions to capability monotonicity are facilities for exception handling and compartmentalization using sealed capabilities, which allow non-monotonicity in a controlled manner to enable software to gain access to additional data capabilities.
@@
The bounds information stored together with the virtual address forms the basis for the memory-safety properties provided by CHERI. Each allocation made by a program running on a CHERI-capable processor is associated with a capability that describes, in addition to the address, the valid bounds of the object (or sub-object) in memory. This empowers CHERI hardware to provide inherent spatial memory-safety properties. Extensions to the CHERI software architecture have explored adding temporal-safety properties to heap-based allocations [86, 82, 25].
@@
III Design
As a motivating example, consider the uninitialized pointer-dereference vulnerability in the Linux kernel shown in LABEL:lst:example. This vulnerability allows control-flow hijacking due to the backlog pointer ❶ being dereferenced in an uninitialized state ❹ if cp->eng_st != ENGINE_IDLE ❷. An attacker could exploit this vulnerability to achieve arbitrary code execution by spraying the stack to control the value of backlog and make it point to attacker-controlled code [47].
Our goal is to detect the first use of the uninitialized pointer in the if-clause already at ❸, unlike the current methods that automatically zero out memory [50, 7]. Our conditional permissions (CPs) offer the same protection against uninitialized memory use as automatic initialization but without the run-time overhead. To achieve this, our implementation of conditional capabilities relies on a novel hardware extension to the CHERI capability architecture we call “Mon CHÈRI”.
Conditional capabilities add, similar to capability sealing (see § II-C), a controlled way of introducing non-monotonicity to the CHERI design. They allow software to temporarily drop permissions, which are then implicitly regained as memory operations fulfilling the condition of the associated CP are met. The design, as we explain in § III-A, ensures that conditional capabilities are guaranteed not to elevate privileges beyond those granted by the corresponding conventional capability.
@@ III Design
Listing 1: An uninitialized pointer dereference vulnerability in the Linux kernel patched in April 2015 [36]. The listing shows all references to the backlog variable in the queue_manag() function from drivers/crypto/mv_cesa.c.
@@
(a) Write-before-Read policy enforced memory area
(b) After occurrence of a store between LB and OB
Figure 2: Write-before-Read capability life cycle: (a) freshly allocated memory under the Write-before-Read policy (blue) is writable only and becomes partially readable (b) after a store operation. UB: upper bound, LB: lower bound, OB: operation bound
@@
(a) High-level overview of the conditional capability-enhanced LLVM compiler. Circled letters indicate the different usage models for conditional capabilities: A Write-before-Read compiler option, B function-level annotations, and C variable-level annotations. Components with a darker background and circled numbers indicate additions compared to the conventional CHERI-enhanced LLVM compiler.
@@
Figure 3: High-level system architecture for our prototype conditional capability-enhanced LLVM compiler and CHERI processor.
III-A Conditional capabilities
@@
Conditional permissions
To describe conditional permissions, we denote conventional permissions, that remain unchanged throughout the lifetime of a capability, as $\mathcal{P^{U}}$, operations on memory as $\mathcal{X}$, and CPs as $\mathcal{P^{C}}$. A $\mathcal{P^{C}}$ is granted on the condition that $\mathcal{X}$ occurs: $\mathcal{X}\implies\mathcal{P^{C}}$ (if $\mathcal{X}$, then $\mathcal{P^{C}}$). If $\mathcal{X}$ does not occur, $\mathcal{P^{C}}$ is not granted. On an architectural level, operations ($\mathcal{X}$) are limited to load, store and execute, while permissions ($\mathcal{P^{U}}$ and $\mathcal{P^{C}}$) can be read, write and execute. For simplicity we will use the terms Write, Read, and Execute when referring to both $\mathcal{X}$ and $\mathcal{P^{C}}$ in CP names. The effective permissions ($\mathcal{P}$) are a subset of the conventional and the conditional permissions:
$\mathcal{P}=\begin{cases}\mathcal{P^{U}}&if\text{ }\mathcal{P^{C}}=\emptyset\\ \mathcal{P^{U}}\land\left(\mathcal{X}\implies\mathcal{P^{C}}\right)&if\text{ }\mathcal{P^{C}}\neq\emptyset\\ \end{cases}$
Operation-specific bounds
Monitoring the memory range for which $\mathcal{P^{C}}:\mathcal{X}\implies\mathcal{P^{C}}$ holds requires tracking operation-specific bounds. Each operation bound tracks the subset of memory within a capability’s conventional bounds for which $\mathcal{X}$ has occurred. To accommodate multiple CPs at the same time, $\mathcal{X}\implies\mathcal{P^{C}}$ and $\mathcal{Y}\implies\mathcal{Q^{C}}$ where $\mathcal{X}\neq\mathcal{Y}$ requires two distinct operation bounds. Conversely, for $\mathcal{X}\implies\mathcal{P^{C}}$ and $\mathcal{X}\implies\mathcal{Q^{C}}$, where $\mathcal{P^{C}}\neq\mathcal{Q^{C}}$, one operation bound is sufficient.
Figure 2 illustrates an example use case for our conditional capabilities: Write-before-Read. At the beginning of the capability’s lifetime (Figure 2.a), the capability refers to a writable memory area with upper and lower bounds. If a store occurs in the memory area (Figure 2.b), the operation bounds increase by the size of the store operand on the hardware level.
The advantage of our design is that it allows capability hardware to enforce a variety of novel access-control permissions beyond conventional RWX. We identify the CPs and corresponding use cases in Table I. This set of CPs allows the definition of “single-use” code and data that enable policies that can be used to harden software against run-time attacks that aim to reuse existing program resource. Such attacks, e.g., code-reuse attacks such as return-oriented-programming [68], can use pre-existing sequences of instructions to craft malicious control flows that can be used to compromise program behavior. Code that a process executes only once, e.g., a fixed sequence of initialization instructions, could, using the Execute-Once CP be invoked once during program initialization and is subsequently rendered useless for code-reuse attacks later in the process’ lifetime.
@@
III-B High-Level System Architecture
Figure 3 depicts the high-level system architecture for our prototype conditional capability-enhanced LLVM compiler and CHERI processor. Inputs and process blocks in a darker color indicate the changes to a conventional CHERI-enabled compiler and processor architecture. For conciseness, we focus on Write-before-Read CPs in the following description.
In our prototype conditional capabilities are exposed to developers through the three alternative usage models: A a Write-before-Read compiler option, B function-level annotations, and C variable-level annotations. Our current design of the compiler option A applies Write-before-Read to all stack variables, but we discuss possibilities for more intelligent heuristics in § VII. Function- and variable-level annotations are implemented as Clang function and variable attributes, respectively (shown in LABEL:lst:attributes), which gives developers control over which variables Write-before-Read is applied to.

### v2

# Mon CHÉRI: Mitigating Uninitialized Memory Access with Conditional Capabilities

@@ Abstract
Up to 10% of memory-safety vulnerabilities in languages like C and C++ stem from uninitialized variables. This work addresses the prevalence and lack of adequate software mitigations for uninitialized memory issues, proposing architectural protections in hardware. Capability-based addressing, such as the University of Cambridge’s CHERI, mitigates many memory defects, including spatial and temporal safety violations at an architectural level. CHERI, however, does not handle undefined behavior from uninitialized variables. We extend the CHERI capability model to include “conditional capabilities”, enabling memory-access policies based on prior operations. This allows enforcement of policies that satisfy memory-safety objectives such as “no reads to memory without at least one prior write” (Write-before-Read). We present our architecture extension, compiler support, and detailed evaluation of our approach on the QEMU full-system simulator and a modified FPGA-based CHERI-RISCV softcore. Our evaluation shows conditional capabilities are practical, with high detection accuracy while adding a small ($\approx$ 3.5%) overhead which is comparable to the cost of baseline CHERI capabilities.
1 Introduction
@@
Uninitialized variables, variables that are declared but not assigned a value, are a well-known source of software defects in the C-family of programming languages. In general, all run-time allocated memory in C, C++, and even in modern languages like Rust, starts out as uninitialized. In this state, the value of the memory is indeterminate and may not reflect a valid state for the variable type. Attempting to interpret uninitialized memory results in undefined behavior, which can cause security vulnerabilities. In particular, uninitialized memory may expose residual data from previously deallocated data structures. This may inadvertently result in information disclosure, ranging from leaked pointer values [1] to the exposure of cryptographic keys [2]. Leaked pointer values can be exploited by attackers to bypass address space layout randomization (ASLR) [1], while the use of uninitialized variables can enable arbitrary code execution attacks [3]. Following recent surveys, use-before-initialized conditions account for a sizable 10% of memory-safety vulnerabilities in the wild [4, 5, 6].
@@
Hardware-assisted defenses against memory-safety issues [7] are motivated by the need to reduce the overhead of run-time defenses through hardware/software co-designs and willingness by processor manufacturers to incorporate mechanisms for software security into their designs [8, 9, 10]. A prominent example is Capability Hardware Enhanced RISC Instructions (CHERI) [11], a joint research project of SRI International and the University of Cambridge. CHERI extends instruction-set architectures (ISAs) with fine-grained memory protection and software compartmentalization.
@@
In its default configuration, CHERI provides spatial safety through capability-based addressing. While this allows software to preclude many classes of memory-safety defects, CHERI does not address uninitialized variables. Microsoft Security Response Center (MSRC) conducted a comprehensive analysis of vulnerabilities reported in 2019 to assess the CHERI ISAv7 [4]. The findings indicate that only 31% of the reported vulnerabilities could have been mitigated through the default configuration of CHERI. An additional 24% could be mitigated by CHERI when configured to provide partial temporal safety under the Cornucopia mechanism [12]. The analysis further highlights that at least 12% of the assessed vulnerabilities could have been mitigated if CHERI would protect against uninitialized access. This work explores and evaluates extensions for CHERI to mitigate those 12%.
@@
In this paper, we extend the CHERI capability model to express memory-access policies that eliminate undefined behavior associated with uninitialized memory. We introduce conditional permissions (CPs) to capability-based addressing. CPs can express memory-access policies that take previous operations on memory into account. This enables conditional capabilities with fine-grained policies that satisfy different instances of memory-safety objectives at different granularities, such as “no reads of memory which has not been the subject of at least one write” (Write-before-Read) or “this memory can be written to only once” (Write-Once). We describe these CPs in § 3. CPs are enforced by introducing the notion of operation-specific bounds to capability-based addressing:
  • We introduce conditional permissions and conditional capabilities for capability-based addressing (§ 3).
@@
  • We integrate conditional capabilities to CHERI-RISC-V in prototypes based on the QEMU full-system emulator and an FPGA softcore based on Flute64Cheri IP (§ 4).
@@
  • We add support for Write-before-Read conditional permissions to the CHERI-enabled Clang/LLVM compiler (§ 4.3) and embedded memory allocators (§ 4.4).
@@
  • We evaluate CPs using $>1000$ NIST Juliet test suite test cases for uninitialized variables and using EEMBC CoreMark performance benchmarks (§ 5).
@@
Our results show that CPs achieve 100% detection with only six false positives ($\approx$ 1%) for the Juliet test suite, where the false positives do exhibit uninitialized (but non-vulnerable) accesses. We further show that CPs on CHERI-RISC-V impose only $\approx$ 3.5% performance overhead in addition to that added by CHERI, compared to benchmarks on an unmodified RISC-V softcore (7% combined). Consequently, CP overhead is comparable to that of CHERI. In summary, our work effectively and efficiently addresses uninitialized memory vulnerabilities, combining very high detection accuracy with performance penalties that are substantially lower than other detection mechanisms in hardware or software. Our conditional capability-enhanced QEMU and toolchain protototypes, and evaluation artifacts are available at https://github.com/conditionalcapabilities.
2 Background
How prevalent are uninitialized memory vulnerabilities?
@@ 2 Background
The prevalence of uninitialized memory as a source of vulnerabilities has been highlighted in statistics based on the Common Vulnerability Enumeration (CVE) program. MSRC reports that, between 2017 and 2019, uninitialized memory vulnerabilities accounted for $\approx$ 5–10% of the Common Vulnerability Enumerations (CVEs) issued by Microsoft[4, 5]. In 2019, uninitialized memory accesses were the fourth largest class of memory defects (10%), after spatial (44%), temporal (29%), and type confusion (14%). Similarly, a more recent survey of memory-safety CVEs between 2015 – 2022 by Sutter [6] indicates that use-before-initialized conditions account for $\approx$ 9% of all reported memory-safety vulnerabilities, after lifetime safety (49%), bounds safety (18%), and type confusion (11%), indicating that uninitialized memory accesses remain an issue in industrial code bases that impacts system security.
2.1 Capability-Based Addressing
@@
Interest in capability-based addressing diminished with the introduction of memory management units (MMUs) that, in addition to performing address translation between virtual and physical memory addresses, also manage access control to virtual memory. However, capabilities differ fundamentally from the access control in MMUs: whereas MMUs associate permissions to individual memory pages, capabilities associate permissions to the references used to address memory.
2.2 The CHERI Capability Architecture
@@
CHERI, which stands for Capability Hardware Enhanced RISC Instructions, is an ISA extension for a capability-based architectural protection model and hardware-software co-design. The CHERI architecture extends an underlying conventional ISA with hardware-supported capabilities that are used to protect virtual addresses used as code or data pointers. The CHERI ISA specification [16] defines the representation of capabilities held in registers and memory, as well as capability-aware instructions to manipulate them. Currently, implementations of CHERI exist for MIPS, RISC-V, and Armv8-A instruction sets. CHERI-enabled processors have been developed by Arm [17], Microsoft [18], as well as in the RISC-V ecosystem.
@@
Figure 1 shows the in-memory representation of a CHERI capability. Each capability is double the width of the native integer pointer type of the baseline architecture: 128 bits on 64-bit platforms and 64 bits on 32-bit platforms. One additional bit, the validity tag ➀, is stored separate from the capability and is protects its integrity: any manipulation of the capability in-memory by non-capability-aware instructions invalidates its tag. Capability-aware instructions maintain the tag as long as certain architectural invariants are met. This prevents direct in-memory manipulations and injection of arbitrary data as capabilities. The permissions ➁ control how the capability can be used and consists of the permissions described in § 2.1. The object type ➂ allows capabilities to be temporarily “sealed”, which renders the capability unusable until it is “unsealed” by a special instruction. Sealing is used by CHERI to implement opaque pointer types and fine-grained in-process isolation. The bounds ➃ describe a lower and upper bound relative to the baseline architecture address ➄, which limits the portion of address space the capability is able to access. To reduce the in-memory footprint of capabilities, the bounds are stored in a compressed format [19] with both bounds in 28 bits (for a 64-bit address), loosing precision as the object size increases.
@@
New capabilities in the CHERI architecture are always derived from an existing capability. The heritage of all capabilities can thus be traced back to the initial capabilities made available to firmware at boot time. CHERI enforces monotonicity on newly created capabilities, ensuring that capabilities constructed by a capability-aware instruction cannot possess permissions or bounds that exceed those of the original capability. The only exceptions to capability monotonicity are facilities for exception handling and compartmentalization using sealed capabilities, which allow non-monotonicity in a controlled manner to enable software to gain access to additional data capabilities.
@@
The bounds information stored together with the virtual address forms the basis for the memory-safety properties provided by CHERI. Each allocation made by a program running on a CHERI-capable processor is associated with a capability that describes, in addition to the address, the valid bounds of the object (or sub-object) in memory. This allows CHERI to provide inherent spatial memory-safety properties. Extensions to the CHERI software stack have explored adding temporal-safety properties to heap-based allocations [20, 12, 21] and sandboxing [22].
@@
3 Conditional Capability Design
Consider the uninitialized pointer-dereference vulnerability in the Linux kernel shown in LABEL:lst:example. This vulnerability allows control-flow hijacking due to the uninitialized backlog pointer ❶ being dereferenced at ❹ when cpg->eng_st != ENGINE_IDLE ❷. An attacker could exploit this vulnerability to achieve arbitrary code execution by spraying the stack to take control of the value of backlog and make it point to attacker-controlled code [3].
Our goal is to detect the first use of the uninitialized pointer in the if-clause at ❸. Unlike methods that automatically zero out memory [23, 5], conditional permissions (CPs) offer the same protection against uninitialized memory use but with smaller and more predictable run-time overhead.
@@ 3 Conditional Capability Design
Listing 1: Excerpt from Linux’ queue_manag() function defined in drivers/crypto/mv_cesa.c showing an uninitialized pointer dereference patched in April 2015 [24].
@@
(a) Initial state of memory area.
(b) State after store between LB and OB
(c) Final state when OB reaches UB.
Figure 2: Write-before-Read conditional capability state transitions: (a) newly allocated memory under the Write-before-Read CP is write-only between the LB and UB and becomes gradually readable (b) after store operations advance the OB. In the final state (c), the OB has reached the UB and the conditional capability behaves identical to a corresponding conventional capability.
@@
(a) Overview of the conditional capability-enhanced LLVM compiler. Circled letters indicate the different usage models for conditional capabilities: A Write-before-Read compiler option, B function-level annotations, and C variable-level annotations. Components with a darker background and circled numbers indicate additions compared to the conventional CHERI-enhanced LLVM compiler.
@@
Figure 3: High-level system architecture of the conditional capability-enhanced LLVM compiler and CHERI processor.
3.1 Challenges
Integrating CPs into capability-based addressing presents several challenges. A key issue is managing the tracking of uninitialized data. Previous approaches [25, 26, 27] assume that data between the lower bound and address pointed to by the capability is initialized, while uninitialized data falls between the address and the upper bound. However, this approach has limitations, as discussed in § 6. To address this, we propose operation-specific bounds (OBs), which precisely track writes. OBs must be updated when a program writes to a region of memory to reflect the new bounds. Otherwise, stale or misaligned bounds can lead to false positives.
Storing and updating operation bounds.
To store and update OBs efficiently, we propose leveraging unused bits in the baseline architecture address (Figure 1). This minimizes changes to the underlying hardware but makes significant compression of the OB necessary. Further details on the hardware modifications are provided in § 3.3 and § 4.
Integration with existing architectures.
To be practical, CPs must integrate seamlessly with existing capability architectures, such as CHERI, that are gaining industry adoption. As a case study, we integrate CPs to CHERI while ensuring full compatibility with the underlying RISC-V ISA. In § 4, we describe the integration process and our hardware extension, “Mon CHÉRI”, which introduces only minor modifications to the CHERI capability representation (§ 4.2).
Maintaining capability monotonicity.
CPs add, similar to capability sealing (§ 2.2), controlled non-monotonicity to the CHERI design. They allow temporary suspension of permissions, which are regained once the associated condition is met. Our design (§ 3.2) guarantees that CPs do not to elevate privileges beyond the original capabilities.
Maintaining capability linearity.
Capability linearity refers to the consistency and integrity of OBs across a program’s execution. To maintain this, we propose a new compiler pass to ensure that capabilities are properly updated as the program executes. This pass is implemented in the LLVM-based Mon CHÉRI toolchain, as detailed in § 3.3 and § 4.3.
Minimizing run-time overhead.
A concern with any memory access control mechanism is the potential for performance degradation. For Mon CHÉRI, we ensure that checks involved in monitoring memory accesses do not add significant latency or computational overhead during normal program execution. In § 4.1, we explain how the OB checks are carefully designed to minimize additional cycles per instruction.
3.2 Conditional Capabilities
@@
Conditional permissions.
To describe conditional permissions, we denote conventional permissions that remain unchanged throughout the lifetime of a capability, as $\mathcal{P^{U}}$, operations on memory as $\mathcal{X}$, and CPs as $\mathcal{P^{C}}$. A $\mathcal{P^{C}}$ is granted on the condition that $\mathcal{X}$ occurs: $\mathcal{X}\implies\mathcal{P^{C}}$ (if $\mathcal{X}$, then $\mathcal{P^{C}}$). If $\mathcal{X}$ does not occur, $\mathcal{P^{C}}$ is not granted. On an architectural level, operations ($\mathcal{X}$) are limited to load, store and execute, while permissions ($\mathcal{P^{U}}$ and $\mathcal{P^{C}}$) can be read, write and execute. For simplicity, we will use the terms Write, Read, and Execute when referring to both $\mathcal{X}$ and $\mathcal{P^{C}}$ in CP names. The effective permissions ($\mathcal{P}$) are a subset of the conventional and the conditional permissions:
$\mathcal{P}=\begin{cases}\mathcal{P^{U}}&if\ \mathcal{P^{C}}=\emptyset\\ \mathcal{P^{U}}\land\left(\mathcal{X}\implies\mathcal{P^{C}}\right)&if\ \mathcal{P^{C}}\neq\emptyset\\ \end{cases}$
Operation-specific bounds.
Tracking the memory range for which $\mathcal{P^{C}}:\mathcal{X}\implies\mathcal{P^{C}}$ requires operation-specific bounds. Each operation bound (OB) tracks the subset of memory within a capability’s conventional bounds for which $\mathcal{X}$ has occurred. To accommodate multiple CPs at the same time, $\mathcal{X}\implies\mathcal{P^{C}}$ and $\mathcal{Y}\implies\mathcal{Q^{C}}$ where $\mathcal{X}\neq\mathcal{Y}$ requires two distinct OBs. Conversely, for $\mathcal{X}\implies\mathcal{P^{C}}$ and $\mathcal{X}\implies\mathcal{Q^{C}}$, where $\mathcal{P^{C}}\neq\mathcal{Q^{C}}$, one OB is sufficient.
Figure 2 illustrates an example use case for our conditional capabilities: Write-before-Read. In its initial state (Figure 2.a), the capability refers to a writable memory area with upper and lower bounds. If a store occurs in the memory area (Figure 2.b), the OB increases by the size of the store operand on the hardware level.
An advantage of our design is that it allows capability hardware to enforce a variety of novel access-control permissions beyond conventional RWX. We identify the CPs and corresponding use cases in Table I. This set of CPs allows the definition of “single-use” code and data that enable policies that can be used to harden software against run-time attacks that aim to reuse existing program resource. Such attacks, e.g., code-reuse attacks such as return-oriented-programming [28], can use pre-existing sequences of instructions to craft malicious control flows that can be used to compromise program behavior. Code that a process executes only once, e.g., a fixed sequence of initialization instructions, could, using the Execute-Once CP be invoked once during program initialization and is subsequently rendered useless for code-reuse attacks later in the process’ lifetime.
@@
3.3 High-Level System Architecture
Figure 3 depicts the high-level system architecture for our prototype conditional capability-enhanced LLVM compiler and CHERI processor. Inputs and process blocks in a darker color indicate the changes to a conventional CHERI-enabled compiler and processor architecture. For conciseness, in this description, we focus on Write-before-Read CPs.
In our prototype, conditional capabilities are exposed to developers through the three alternative usage models: A a Write-before-Read compiler option, B function-level annotations, and C variable-level annotations. Our current design of the compiler option A applies Write-before-Read to all stack variables, but we discuss possibilities for more intelligent heuristics in § 7. Function- and variable-level annotations are implemented as Clang function and variable attributes, respectively (shown in LABEL:lst:attributes), which gives developers control over which variables Write-before-Read is applied to.